OpenStudy (anonymous):

Express the perimeter of the triangle as a polynomial. A. 9x + 8 B. 9x – 8 C. –9x + 6 D. 8x – 6

OpenStudy (anonymous):

@MathLegend

hartnn (hartnn):

perimeter = sum of sides

hartnn (hartnn):

\( 3x+4 ~ + 5x-2~ + x-10 = ...?\)

OpenStudy (anonymous):

ok let me see if i can solve that.

hartnn (hartnn):

combine like terms \(3x+5x+x = (3+5+1)x = ... \\ 4-2-10=...\)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

b

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^